OpenStudy (anonymous):

a body is thrown vertically up with a speed of 10 m/s from the top of a wall of height 3.45m the height from which another body should be dropped simultaneously such that two bodies reach the ground in same time (g=10m/s2) explanation??

OpenStudy (aravindg):

guys pls dont help

OpenStudy (aravindg):

this qn was answered before

OpenStudy (aravindg):

he is posting same qn again

OpenStudy (anonymous):

that was a wrong answer

OpenStudy (aravindg):

no it was ryt

OpenStudy (aravindg):

@experimentX pls dont post an answer

OpenStudy (anonymous):

aravid gave me the answer as 26m but it is not there in the option

OpenStudy (anonymous):

the answer is 24.5m

OpenStudy (aravindg):

take a look at this http://openstudy.com/updates/4fa62981e4b029e9dc35e6e7

OpenStudy (experimentx):

calculate the time of flight!!! while going up v = u + at, u = 10, v = 0, t= find it ... let it be t1 = 1s from prev question also distance v^2 = u^2 + 2as calculate time of while falling s+3.45 = ut + 1/2 gt^2 <--- get t1 from here, u=0 here s = u(t1 + t2) + 1/2 a(t1 + t2)^2, u =0 find s using this formula,

OpenStudy (aravindg):

wow superb answer experiment :P

OpenStudy (experimentx):

 To visualize it propertly